• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 12:42:51 on August20,2003

    Entry number 112094

    Followups:

    keyword=Controlled access for EPR ended (problem not fixed yet). Ready for beam.

    Vince went to the hall to re-align EPR fiber. It had little effect. So probably something else was not working: photodiode or fiber. Will need a longer access to fix it. Controoled access ended at 13:30. I moved target to Empty. We are ready for beam.